OpenStudy (anonymous):

Limit X-->0 csc(3x)/cot(4x) ?

OpenStudy (abb0t):

try changing everything to sines and cosines \[\csc = \frac{ 1 }{ \sin }, \cot = \frac{ \cos }{ \sin }\]